Peter is 7 years older than his wife Wendy. The sum of their ages is 63 find Peter's age.
Answer by ikleyn(52786)

.

            I will show you TWO WAYS to solve the problem.


1-st way

If P is the Peter's age, then (P-7) is the Wendy's age.


From the condition,


    P + (P-7) = 63,

    2P = 63 + 7 = 70,

     P =        = 70/2 = 35.


ANSWER.  Peter is 35 years old.


2-nd way

From the condition, you have these two equations

    
    P + W = 63

    P - W =  7


Add the equations and see what will happen.